# Subsystem Information Capacity in Random Circuits and Hamiltonian Dynamics

### v2

@@ III.2.1 Random circuit with one-to-one encoding
We find that the steady mutual information for the random unitary circuit without mid-circuit measurements shows a step function behavior: for $L_{A}<L/2$, $MI=0$ while for $L_{A}>L/2$, $MI=2$ (pattern (d) in Fig. 2). In other words, the information can be perfectly decoded from a subsystem with more than half the system while the information is totally lost for a subsystem of less than half of the system. This result indicates full information scrambling and is consistent with the philosophy of Hayden-Preskill thought experiment Hayden and Preskill (2007). Considering the information conservation law $I(A:R)+I(\bar{A}:R)=2$, the steady-state mutual information applies to any subregion of a given size no matter whether $A$ has included the initial $E$ qubit. This is understandable since the information is fully scrambled in structureless random quantum circuits rendering the initial position of the encoding irrelevant. It is worth noting that this typical step function behavior in the steady-state mutual information is a universal feature of thermalized systems, which we will revisit in the context of fermionic Hamiltonian dynamics.
@@
Figure 5: SIC dynamics in random quantum circuit with different subsystem sizes $L_{A}$ using one-to-one encoding scheme. System size $L=512$, average over 1000 configurations. Measurement probability $p_{m}=0$. The dashed line corresponds to the analytical results from the effective statistical model.
@@
Next, we investigate the dynamics of $MI(x,t)$ with varying $t$ before reaching steady states. Fig. 5 illustrates the information dynamics for random quantum circuits without mid-circuit measurements. In this figure, we also present the predictions from the analytical framework using the entanglement membrane picture and find that the results match remarkably well with the theoretical predictions. To gain a deeper understanding of the dynamics, , we show the dominant entanglement membrane picture for different times in Fig. 6. In each spin configuration plot, the time direction is oriented upwards. The upper panel in each case shows the dominant configuration for $S_{A}$ with top boundary conditions $\mathbb{C}$ (purple color) in the region $A$ and $\mathbb{I}$ on the region $\bar{A}$ as well as the reference qubit $R$. The lower panel in each case shows the dominant configuration for $S_{AR}$, and the only difference compared to the effective model for $S_{A}$ is that the reference qubit is applied with the boundary condition $\mathbb{C}$. To compute the free energy contribution of each given configuration, we compute the contribution from the tension energy of the domain walls and the overlaps from the bottom boundary.
@@
For example, for $L_{A}<L/2$, the top-boundary triangular domain wall is only geometrically possible for $v_{E}t>L_{A}/2$ and contributes to the free energy as $L_{A}$ as we explained in Sec. III.1. Similarly, the vertical domain wall contributes free energy as $v_{E}t$, where $t$ is the domain wall length. As the vertical domain wall and the triangular domain wall near the top boundaries compete, the phase transition occurs for the time $t=L_{A}/(2v_{E})$. To compute the entropy, we have to also consider the contribution from the bottom boundary overlap. Since the input state beyond the entangled qubit is in the form of a product state, the bottom boundary beyond the entangled qubit is all free, in the sense that for each qubit $i$ we have $\text{Tr}(\mathbb{C}_{i}\rho_{i}^{\otimes r})=\text{Tr}(\mathbb{I}_{i}\rho_{i}^{\otimes r})=1$. For the entangled qubit with reference qubit $E\cup R$, we have $\text{Tr}(\mathbb{C}_{E}\mathbb{C}_{R}\rho_{B}^{\otimes r})=\text{Tr}(\mathbb{I}_{E}\mathbb{I}_{R}\rho_{B})=1$ and $\text{Tr}(\mathbb{C}_{E}\mathbb{I}_{R}\rho_{B}^{\otimes r})=\text{Tr}(\mathbb{I}_{E}\mathbb{C}_{R}\rho_{B}^{\otimes r})=d^{-r}$, where $\rho_{B}=\frac{1}{2}(|0_{E}0_{R}\rangle+|1_{E}1_{R}\rangle)(\langle 0_{E}0_{E}|+\langle 1_{E}1_{R}|)$ is the initial Bell pair. In sum, for $L_{A}<L/2$ and $v_{E}t<L_{A}/2$, we have $S_{A}=2v_{E}t+1$. Similarly, we have $S_{AR}=2v_{E}t$ which lead to the mutual information $I(A:R)=S_{A}+1-S_{AR}=2$. The physical meaning is that the information is fully kept in the subsystem at short times $t<L_{A}/(2v_{E})$. On the other hand, the top-boundary triangular weight will dominate, and gives $S_{A}=L_{A}$ and $S_{AR}=L_{A}+1$ at late times, leaving the mutual information $I(A:R)=0$. The prediction is consistent with numerical results in Fig. 5. We find that for subsystems smaller than half the system, the mutual information first is 2 and at some later time $t^{*}$ suddenly drops to zero. The timescale $t^{*}$ for the mutual information drop is exactly given by $L_{A}/(2v_{E})$, and this setting explains why $v_{E}$ can be regarded as a velocity.
@@
The analysis utilizing the effective statistical model can be extended to other information metrics, such as Holevo information mentioned in Sec. II.3. In this case, the upper panel for each time slice and $L_{A}$ are kept unchanged to account for the first term in Holevo information, $S_{A}$. The second term for Holevo information is characterized by the statistical model with free boundary conditions on the entangled qubit $E$. For small $t$, the entanglement membrane picture remains the same as $S_{AR}$, yielding the same plateau dynamics starting from $\chi_{A}=1$. For steady-state Holevo information, the second term is $L_{A}$ for $L_{A}\leq L/2$ and $L-L_{A}$ for $L_{A}>L/2$. The steady-state $\chi_{A}(t\rightarrow\infty)$ is thus 00 and $1$ for $L_{A}\leq L/2$ and $L_{A}>L/2$, respectively. Remarkably, the results for mutual information and Holevo information show qualitative difference when $L_{A}=L/2$: while the steady-state mutual information is half as the initial value, the steady-state Holevo information is zero.
Figure 7: SIC dynamics in random quantum circuit with different subsystem sizes $L_{A}$ using one-to-one encoding scheme with mid-circuit measurement probability $p_{m}=0.1$. (a) $L=512$ and (b) $L=64$, the results are averaged over 1000 random circuit configurations.
@@
With the mid-circuit measurement $p_{m}>0$, the information dynamics is shown in Fig. 7. For large-size systems, the information decay is categorized into two stages. In the early stage, the information stored in the subsystem universally drops due to the measurements. Since we are in the noise-resilient phase of MIPT, the final steady information is non-zero. The time scale $t^{*}$ to approach the measurement-induced steady state is rather small in the order of $O(10)$ for $p_{m}=0.1$ and agnostic to the system size. For any subsystem size $L_{A}>2v_{E}t^{*}$, the corresponding mutual information with the reference qubit $R$ will experience two stages of dynamics separately. At the first stage, the information drops due to the monitored dynamics and stays at the plateau induced by measurements for $t^{*}<t<L_{A}/(2v_{E})$. At the second stage with $t>L_{A}/(2v_{E})$, the information begins to drop again from the universal plateau due to the information scrambling to the complementary system $\bar{A}$. For small systems where $L_{A}\sim 2v_{E}t^{*}$, the two stages of information dynamics become mixed as shown in Fig. 7(b), i.e. the information is spreading to the subsystem $\bar{A}$ and leak to the environment at the same time.
III.2.2 Random Floquet circuit with one-to-one encoding
@@
In this subsection, we investigate the information dynamics in random Floquet circuits, i.e. random circuits with time periodicity. These circuits exhibit distinct dynamical phases depending on whether the local random gates are sampled from a random Haar ensemble or a random Clifford ensemble. The former ensemble leads to thermal phases Sünderhauf et al. (2018) while the latter ensemble leads to Anderson localization phases in one dimension Farshi et al. (2022) and thermal phases in higher dimensions Farshi et al. (2023). To facilitate the large-scale numerical simulation, we focus on the 1D random Clifford Floquet circuits as the testbed. This model can reveal the information dynamics in Anderson localization systems while other chaotic Floquet systems are believed to show similar full information scrambling dynamics as discussed in Sec. III.2.1.
@@
Fig. 8 shows the steady-state mutual information with respect to different subsystem sizes $L_{A}$. In Anderson localization phases, the mutual information curve converged in thermodynamic limit with $L_{A}$, indicating a system size irrelevant localization length. The curve will be always at value $2$ if the x-axis is $L_{A}/L$ in the thermodynamic limit. In other words, the localization length restricts the capacity of information scrambling, and no information can spread too far away beyond the localization length. Therefore, the output subsystem with size on the order of localization length is sufficient to extract most of the initial information stored at the beginning.
III.2.3 Random circuit with one-to-all encoding
@@
Fig. 9 presents steady-state mutual information with different subsystem sizes $L_{A}$. We find the result is consistent with one-to-one encoding. Again, the results only rely on the subsystem size and are independent of the subsystem’s position which is unsurprising as there is no specific entangled qubit at all due to the global nature of the one-to-all encoding scheme. The corresponding information dynamics is shown in Fig. 10. The initial value for $MI$ is $1$ for different system sizes as expected. For subsystems with size $L_{A}>L/2$ ($L_{A}<L/2$), the information capacity increases (decreases) with the evolved time from $1$ and vice versa. The entanglement membrane picture, as shown in Fig. 11, offers a comprehensive explanation for the steady and dynamical behaviors of subsystem information capacity. It is worth noting that the nonuniform configuration at the bottom boundary gives $1$ extra unit of free energy due to the boundary condition imposed by GHZ states. The timescale at which the mutual information begins to deviate from the initial value $1$ is also determined by $L_{A}/(2v_{E})$ for $L_{A}<L/2$, which is consistent with the numerical results.
@@
III.2.4 Random Floquet circuit with one-to-all encoding
Since the initial values of $MI$ are $1$ for all subsystems with one-to-all encoding scheme, an interesting question arises on the fate of such information dynamics in Anderson localization systems. We here report the information dynamics results with one-to-all encoding in random Clifford Floquet circuits. The late-time steady-state mutual information curve is depicted in Fig. 12. With increasing system sizes, the steady-state mutual information also converges to $1$ irrespective of the subsystem size $L_{A}<L$. Thus we conclude in the thermodynamic limit, the steady-state information distribution gives $1$ for any $L_{A}<L$ ($MI=2$ for $L_{A}=L$). This fact suggests Anderson localization can stop the information spreading, a phenomenon that becomes more pronounced as shown in Fig. 13. The mutual information for any subsystem size remains essentially unchanged, with minor fluctuations attributed to finite-size effects. In the thermodynamic limit, mutual information is expected to stabilize at $1$ for all times and across all subsystem sizes $L_{A}<L$. The global encoding strategy thus offers a clearer insight into the intrinsic characteristics of Anderson localization, where information spreading is effectively halted.
@@ III.2.4 Random Floquet circuit with one-to-all encoding
III.2.5 Random circuit with many-to-many encoding
@@
Figure 15: SIC dynamics in 1D random quantum circuit with different subsystem sizes $L_{A}$ under finite rate encoding scheme which encodes $L/2$ Bell pairs with $L/2$ ancilla qubits at the beginning. System size is $L=512$ with $p_{m}=0$. The results are averaged over 200 configurations. The information decay is very slow compared to dynamics in the one-to-one encoding case. The dashed line is the analytical result predicted by the entanglement membrane picture.
Figure 16: SIC dynamics in 1D random quantum circuit with different subsystem sizes $L_{A}$ under finite rate encoding scheme which encodes $L/2$ Bell pairs with $L/2$ ancilla qubits at the beginning. System size is $L=512$ with $p_{m}=0.05$. The results are averaged over 200 configurations.
@@
The three stages of the information dynamics, predicted by the dashed line in Fig. 15, can also be perfectly explained via the entanglement membrane picture as shown in Fig. 17. Unlike one-to-one and one-to-all encoding cases, now the bottom boundary condition can have extended overhead. If the $L/2$ ancilla qubits are imposed with boundary condition $\mathbb{C}$, then each configuration $\mathbb{I}$ in the region $E$ will contribute $1$ to the free energy. The difference in the bottom boundary condition induces the emergent stage where the information shows linear decay instead of the sudden drop observed and predicted in previous cases. This new stage is specifically tied to the second column in Fig. 17(a) when $(L_{A}-L/2)/(2v_{E})<t<(L-L_{A})/(2v_{E})$. For $S_{A}$ in this case, the free energy is given as $2\mathbb{E}(v)t+L_{A}-2vt$, the minimization of this term gives $\mathbb{E}^{\prime}(v)=1$, which determines the velocity of the optimal entanglement membrane as $v_{b}$ and thus $S_{A}=L_{A}$ independent of the time $t$. However, $S_{AR}=2v_{E}t$ in this case relies on $t$. Therefore, the target SIC, as the differences between the two terms, has linear $t$ dependence as shown from numerical simulation. For the cases with mid-circuit measurements, we again observe the two stages of information drop: the first stage is universal as information leakage induced by measurements and the second stage resembles the behaviors for $p_{m}=0$ limit which is induced by information propagation.
III.2.6 Random Floquet circuit with many-to-many encoding

### v3

@@ 3.2.1 Random circuit with one-to-one encoding
We find that the steady mutual information for the random unitary circuit without mid-circuit measurements shows a step function behavior: for $L_{A}<L/2$, $MI=0$ while for $L_{A}>L/2$, $MI=2$ (pattern (d) in Fig. 2). In other words, the information can be perfectly decoded from a subsystem with more than half the system while the information is totally lost for a subsystem of less than half of the system. This result indicates full information scrambling and is consistent with the philosophy of Hayden-Preskill thought experiment [7]. Considering the information conservation law $I(A:R)+I(\bar{A}:R)=2$, the steady-state mutual information applies to any subregion of a given size no matter whether $A$ has included the initial $E$ qubit. This is understandable since the information is fully scrambled in structureless random quantum circuits rendering the initial position of the encoding irrelevant. It is worth noting that this typical step function behavior in the steady-state mutual information is a universal feature of thermalized systems, which we will revisit in the context of fermionic Hamiltonian dynamics.
@@
Figure 5: SIC dynamics in random quantum circuit with different subsystem sizes $L_{A}$ using one-to-one encoding scheme with or without mid-circuit measurements. (a) System size $L=512$, mid-circuit measurements $p_{m}=0$, (b) $L=512$, $p_{m}=0.1$, (c) $L=64$, $p_{m}=0.1$, average over 1000 random circuit configurations.
@@
Next, we investigate the dynamics of $MI(x,t)$ with varying $t$ before reaching steady states. Fig. 5(a) illustrates the information dynamics for random quantum circuits without mid-circuit measurements. In this figure, we also present the predictions from the analytical framework using the entanglement membrane picture and find that the results match remarkably well with the theoretical predictions. To gain a deeper understanding of the dynamics, we show the dominant entanglement membrane picture for different times in Fig. 6. In each spin configuration plot, the time direction is oriented upwards. The upper panel in each case shows the dominant configuration for $S_{A}$ with top boundary conditions $\mathbb{C}$ (purple color) in the region $A$ and $\mathbb{I}$ on the region $\bar{A}$ as well as the reference qubit $R$. The lower panel in each case shows the dominant configuration for $S_{AR}$, and the only difference compared to the effective model for $S_{A}$ is that the reference qubit is applied with the boundary condition $\mathbb{C}$. To compute the free energy contribution of each given configuration, we compute the contribution from the tension energy of the domain walls and the overlaps from the bottom boundary.
@@
For example, for $L_{A}<L/2$, the top-boundary triangular domain wall is only geometrically possible for $v_{E}t>L_{A}/2$ and contributes to the free energy as $L_{A}$ as we explained in Sec. 3.1. Similarly, the vertical domain wall contributes free energy as $v_{E}t$, where $t$ is the domain wall length. As the vertical domain wall and the triangular domain wall near the top boundaries compete, the phase transition occurs for the time $t=L_{A}/(2v_{E})$. To compute the entropy, we have to also consider the contribution from the bottom boundary overlap. Since the input state beyond the entangled qubit is in the form of a product state, the bottom boundary beyond the entangled qubit is all free, in the sense that for each qubit $i$ we have $\text{Tr}(\mathbb{C}_{i}\rho_{i}^{\otimes r})=\text{Tr}(\mathbb{I}_{i}\rho_{i}^{\otimes r})=1$. For the entangled qubit with reference qubit $E\cup R$, we have $\text{Tr}(\mathbb{C}_{E}\mathbb{C}_{R}\rho_{B}^{\otimes r})=\text{Tr}(\mathbb{I}_{E}\mathbb{I}_{R}\rho_{B})=1$ and $\text{Tr}(\mathbb{C}_{E}\mathbb{I}_{R}\rho_{B}^{\otimes r})=\text{Tr}(\mathbb{I}_{E}\mathbb{C}_{R}\rho_{B}^{\otimes r})=d^{-r}$, where $\rho_{B}=\frac{1}{2}(|0_{E}0_{R}\rangle+|1_{E}1_{R}\rangle)(\langle 0_{E}0_{E}|+\langle 1_{E}1_{R}|)$ is the initial Bell pair. In sum, for $L_{A}<L/2$ and $v_{E}t<L_{A}/2$, we have $S_{A}=2v_{E}t+1$. Similarly, we have $S_{AR}=2v_{E}t$ which lead to the mutual information $I(A:R)=S_{A}+1-S_{AR}=2$. The physical meaning is that the information is fully kept in the subsystem at short times $t<L_{A}/(2v_{E})$. On the other hand, the top-boundary triangular weight will dominate, and gives $S_{A}=L_{A}$ and $S_{AR}=L_{A}+1$ at late times, leaving the mutual information $I(A:R)=0$. The prediction is consistent with numerical results in Fig. 5(a). We find that for subsystems smaller than half the system, the mutual information first is 2 and at some later time $t^{*}$ suddenly drops to zero. The timescale $t^{*}$ for the mutual information drop is exactly given by $L_{A}/(2v_{E})$, and this setting explains why $v_{E}$ can be regarded as a velocity.
@@
The analysis utilizing the effective statistical model can be extended to other information metrics, such as Holevo information mentioned in Sec. 2.3. In this case, the upper panel for each time slice and $L_{A}$ are kept unchanged to account for the first term in Holevo information, $S_{A}$. The second term for Holevo information is characterized by the statistical model with free boundary conditions on the entangled qubit $E$. For small $t$, the entanglement membrane picture remains the same as $S_{AR}$, yielding the same plateau dynamics starting from $\chi_{A}=1$. For steady-state Holevo information, the second term is $L_{A}$ for $L_{A}\leq L/2$ and $L-L_{A}$ for $L_{A}>L/2$. The steady-state $\chi_{A}(t\rightarrow\infty)$ is thus 00 and $1$ for $L_{A}\leq L/2$ and $L_{A}>L/2$, respectively. Remarkably, the results for mutual information and Holevo information show qualitative difference when $L_{A}=L/2$: while the steady-state mutual information is half as the initial value, the steady-state Holevo information is zero.
@@
With the mid-circuit measurement $p_{m}>0$, the information dynamics is shown in Fig. 5(b)(c). For large-size systems, the information decay is categorized into two stages. In the early stage, the information stored in the subsystem universally drops due to the measurements. Since we are in the noise-resilient phase of MIPT, the final steady information is non-zero. The time scale $t^{*}$ to approach the measurement-induced steady state is rather small in the order of $O(10)$ for $p_{m}=0.1$ and agnostic to the system size. For any subsystem size $L_{A}>2v_{E}t^{*}$, the corresponding mutual information with the reference qubit $R$ will experience two stages of dynamics separately. At the first stage, the information drops due to the monitored dynamics and stays at the plateau induced by measurements for $t^{*}<t<L_{A}/(2v_{E})$. At the second stage with $t>L_{A}/(2v_{E})$, the information begins to drop again from the universal plateau due to the information scrambling to the complementary system $\bar{A}$. For small systems where $L_{A}\sim 2v_{E}t^{*}$, the two stages of information dynamics become mixed as shown in Fig. 5(c), i.e. the information is spreading to the subsystem $\bar{A}$ and leaks to the environment at the same time.
3.2.2 Random Floquet circuit with one-to-one encoding
@@
In this subsection, we investigate the information dynamics in random Floquet circuits, i.e. random circuits with time periodicity. These circuits exhibit distinct dynamical phases depending on whether the local random gates are sampled from a random Haar ensemble or a random Clifford ensemble. The former ensemble leads to thermal phases [42] while the latter ensemble leads to Anderson localization phases in one dimension [40] and thermal phases in higher dimensions [41]. To facilitate the large-scale numerical simulation, we focus on the 1D random Clifford Floquet circuits as the testbed. This model can reveal the information dynamics in Anderson localization systems while other chaotic Floquet systems are believed to show similar full information scrambling dynamics as discussed in Sec. 3.2.1.
@@
Fig. 7 shows the steady-state mutual information with respect to different subsystem sizes $L_{A}$. In Anderson localization phases, the mutual information curve converges in thermodynamic limit with $L_{A}$ as x-axis, indicating a system size irrelevant localization length. The curve will be always at value $2$ if the x-axis is $L_{A}/L$ in the thermodynamic limit. In other words, the localization length restricts the capacity of information scrambling, and no information can spread far away beyond the localization length. Therefore, the output subsystem with size on the order of localization length is sufficient to extract most of information stored at the beginning.
3.2.3 Random circuit with one-to-all encoding
@@
Fig. 8 presents steady-state mutual information with different subsystem sizes $L_{A}$. We find the result is consistent with one-to-one encoding and the curves of different sizes collapse with rescaled subsystem size $L_{A}/L$, contrary to Fig. 7, where late-time curves could collapse with respect to subsystem sizes $L_{A}$. Again, the results only rely on the subsystem size and are independent of the subsystem’s position which is unsurprising as there is no specific entangled qubit at all due to the global nature of the one-to-all encoding scheme. The corresponding information dynamics is shown in Fig. 9. The initial value for $MI$ is $1$ for different system sizes as expected. For subsystems with size $L_{A}>L/2$ ($L_{A}<L/2$), the information capacity increases (decreases) with the evolved time from $1$ and vice versa. The entanglement membrane picture, as shown in Fig. 10, offers a comprehensive explanation for the steady and dynamical behaviors of subsystem information capacity. It is worth noting that the nonuniform configuration at the bottom boundary gives $1$ extra unit of free energy due to the bottom boundary condition imposed by GHZ states. The timescale at which the mutual information begins to deviate from the initial value $1$ is also determined by $L_{A}/(2v_{E})$ for $L_{A}<L/2$, which is consistent with the numerical results.
@@
3.2.4 Random Floquet circuit with one-to-all encoding
Since the initial values of $MI$ are $1$ for all subsystems with one-to-all encoding scheme, an interesting question arises on the fate of such information dynamics in Anderson localization systems. We here report the information dynamics results with one-to-all encoding in random Clifford Floquet circuits. The late-time steady-state mutual information curve is depicted in Fig. 11. With increasing system sizes, the steady-state mutual information also converges to $1$ irrespective of the subsystem size $L_{A}<L$. Thus we conclude in the thermodynamic limit, the steady-state information capacity is always $1$ for any $L_{A}<L$ ($MI=2$ for $L_{A}=L$). This fact suggests that Anderson localization can fully stop the information spreading, a phenomenon that becomes more pronounced as shown in Fig. 12. The mutual information for any subsystem size remains essentially unchanged, with minor fluctuations attributed to finite-size effects. In the thermodynamic limit, mutual information is expected to stabilize at $1$ for all times and across all subsystem sizes $L_{A}<L$. The global encoding strategy thus offers a clearer insight into the intrinsic characteristics of Anderson localization, where information spreading is effectively halted.
@@ 3.2.4 Random Floquet circuit with one-to-all encoding
3.2.5 Random circuit with many-to-many encoding
@@
Figure 14: SIC dynamics in 1D random quantum circuit with different subsystem sizes $L_{A}$ under finite rate encoding scheme which encodes $L/2$ Bell pairs with $L/2$ ancilla qubits at the beginning. System size is $L=512$ with $p_{m}=0$ (a) $p_{m}=0.05$ (b) and $p_{m}=0.1$ (c). The results are averaged over 200 circuit configurations. The information decay is very slow compared to dynamics in the one-to-one encoding case. The dashed line in (a) is the analytical result predicted by the entanglement membrane picture.
@@
The three stages of the information dynamics, predicted by the dashed line in Fig. 14(a), can also be perfectly explained via the entanglement membrane picture as shown in Fig. 15. Unlike one-to-one and one-to-all encoding cases, now the bottom boundary condition can have extended overhead. If the $L/2$ ancilla qubits are imposed with boundary condition $\mathbb{C}$, then each configuration $\mathbb{I}$ in the region $E$ will contribute $1$ to the free energy. The difference in the bottom boundary condition induces the emergent stage where the information shows linear decay instead of the sudden drop observed and predicted in previous cases. This new stage is specifically tied to the second column in Fig. 15(a) when $(L_{A}-L/2)/(2v_{E})<t<(L-L_{A})/(2v_{E})$. For $S_{A}$ in this case, the free energy is given as $2\mathbb{E}(v)t+L_{A}-2vt$, the minimization of this term gives $\mathbb{E}^{\prime}(v)=1$, which determines the velocity of the optimal entanglement membrane as $v_{b}$ and thus $S_{A}=L_{A}$ independent of the time $t$. However, $S_{AR}=2v_{E}t$ in this case relies on $t$. Therefore, the target SIC, as the differences between the two terms, has linear $t$ dependence as shown from numerical simulation. For the cases with mid-circuit measurements, we again observe the two stages of information drop: the first stage is universal as information leakage induced by measurements and the second stage resembles the behaviors for $p_{m}=0$ limit which is induced by information propagation.
3.2.6 Random Floquet circuit with many-to-many encoding